### v1

# A Semantic Invariant Robust Watermark for Large Language Models

Abstract
Watermark algorithms for large language models (LLMs) have achieved extremely high accuracy in detecting text generated by LLMs. Such algorithms typically involve adding extra watermark logits to the LLM’s logits at each generation step. However, prior algorithms face a trade-off between attack robustness and security robustness. This is because the watermark logits for a token are determined by a certain number of preceding tokens; a small number leads to low security robustness, while a large number results in insufficient attack robustness. In this work, we propose a semantic invariant watermarking method for LLMs that provides both attack robustness and security robustness. The watermark logits in our work are determined by the semantics of all preceding tokens. Specifically, we utilize another embedding LLM to generate semantic embeddings for all preceding tokens, and then these semantic embeddings are transformed into the watermark logits through our trained watermark model. Subsequent analyses and experiments demonstrated the attack robustness of our method in semantically invariant settings: synonym substitution and text paraphrasing settings. Finally, we also show that our watermark possesses adequate security robustness. Our code and data are available at https://github.com/THU-BPM/Robust_Watermark.
1 Introduction
As the quality of text generated by large language models (LLMs) continues to improve, it addresses a multitude of practical challenges on one hand, while simultaneously giving rise to a spectrum of new issues on the other. Specifically, the proliferation of LLM-generated text on the Internet may lead to an influx of rumor-based content and text copyright concerns (Rillig et al., 2023). Therefore, the detection and labeling of machine-generated text have become extremely important.
Text watermarking techniques for LLMs usually embed specific information during text generation to allow high-accuracy detection of LLM-generated text. The mainstream approach for embedding such information is to add extra watermark logits on top of the logits generated by the LLM. For example, Kirchenbauer et al. (2023) divide the vocabulary into red and green lists and increase the scores for the green tokens as the watermark logits. However, current watermarking algorithms cannot possess both attack robustness (robustness to modifications of the watermarked text) and security robustness (robustness against watermark cracking). For example, Zhao et al. (2023) demonstrates that global fixed watermark logits enhance attack robustness, yet they compromise security due to vulnerability in word frequency analysis (Sadasivan et al., 2023). Essentially, in current watermark algorithms Kirchenbauer et al. (2023); Zhao et al. (2023); Wang et al. (2023), the watermark logits for each token depend on its preceding tokens. As the number of required preceding tokens increases, watermarking complexity rises, leading to reduced attack robustness but increased security robustness.
To resolve the aforementioned trade-off, we propose a semantic invariant watermarking algorithm that achieves reasonable attack and security robustness. The core motivation is generating watermark logits for each token based on the preceding tokens’ semantics rather than their token IDs. Thus, semantically invariant text modifications do not alter the watermark logits, while the diversity of text semantics increases watermark complexity and guarantees security against watermark cracking. Specifically, to extract semantically invariant features, we use an auxiliary LLM encoder (e.g., BERT) to extract semantic embeddings of the preceding tokens. We then train a small watermark network to transform these embeddings into corresponding watermark logits. The training objective of the watermark network is to ensure a high correlation between the similarities of the output watermark logits and input text embeddings. Also, it’s imperative that the watermark logits exhibit sufficient diversity and are unbiased for each token. To achieve these goals, two training objectives are adopted: a similarity loss and a normalization loss. For the similarity loss, to ensure the diversity of the watermark logits, we first rescale the similarity values between text embeddings to range from -1 to 1, and then make the similarity of the generated watermark logits fit this similarity. To ensure unbiased token selection and achieve bimodal scores of the watermark logits, our normalization loss centers the mean of each row and column within a batch of watermark logits to zero, while making the absolute value of each entry as close as possible. We detect watermarks by averaging the watermark logits across all tokens. If the average is significantly above zero, a watermark is present.
Figure 1: An illustration of our semantic invariant robust watermarking method. Text is input into a generative LLM for token logits and an embedding LLM for text embedding. The embedding is converted into watermark logits via the Watermark Model. LLM logits and watermark logits are then combined for final logits, which decode the next token using any method.
In the experiment, we evaluate the attack robustness of our watermarking algorithm against various semantically invariant perturbations, including text paraphrasing and synonym replacement. Overall, our watermark robustness is comparable to KGW-1 (global watermark logits), which is close to the robustness upper bound achievable by watermark logit-based methods. Additionally, employing the spoofing attack paradigm used in Sadasivan et al. (2023), we evaluate the decryption accuracy of various watermarking methodologies to gauge security robustness. Our algorithm demonstrates favorable security robustness metrics, effectively resolving the previously encountered trade-off between attack and security robustness. Importantly, the watermark logits could be generated in parallel with the LLM logits, resulting in only a marginal latency increase during text generation.
In summary, the contributions of this work are as follows:
  • We propose the first semantically invariant robust watermarking algorithm, which effectively detects watermarks under various semantically invariant perturbations.
  • Our algorithm successfully navigates the trade-off between attack robustness and security robustness that plagued previous methods, achieving high performance in both dimensions.
  • We innovatively propose a watermark model that adeptly transforms semantic embeddings into watermark logits.
2 Related work
Text watermarking technology embeds imperceptible patterns within textual content and employs specialized detection techniques to ascertain the presence of a watermark (Abdelnabi & Fritz, 2021; Yang et al., 2022). Currently, there are two approaches to watermarking text: post-processing after text generation, and incorporating watermarks during the text generation of LLM.
Post-processing methods make minor semantic-preserving alterations to the text, most commonly lexical substitution-based approaches. For instance, Qiang et al. (2023) embedded watermark information by employing a paraphraser-based lexical substitution, ensuring the text’s original semantics remained intact. Yoo et al. (2023) began by selecting stable words to position the watermark and then used a masked language model to generate candidate words, thereby facilitating the embedding and extraction of watermarks. Munyer & Zhong (2023) generated candidate texts through word2vec-based synonym replacements, subsequently comparing semantic similarities among candidates to finalize the watermarked text. Nonetheless, replacing individual words while only considering the unchanged meaning of those words risks degrading overall text quality. In contrast, generative modification methods like Abdelnabi & Fritz (2021) directly generate watermarked text, along with corresponding networks to recover the watermark information. However these methods can still only make minimal changes to the original text, severely limiting the freedom to insert watermarks.
Compared to post-processing methods, integrating watermarks into the text generation process of current generative language models better utilizes their capabilities. Based on the stage at which watermark information is embedded, current watermarking methods for LLMs can be divided into two categories: adding watermarks during token sampling and during logits computation. Christ et al. (2023) proposed a method that embeds watermark information by presetting a random number sequence for token sampling, and detects watermarks by computing correlation between texts and the random number sequence. To improve robustness against text modifications, Kuditipudi et al. (2023) adopted Levenshtein distance for matching texts and the random number sequence, allowing watermarked texts to be detectable even after certain modifications. However, although such token sampling-based watermarking methods minimize the impact on text quality, they also greatly limit the randomness of generated texts, and only work for sampling decoders instead of other decoding methods like beam search. In contrast, another category is to add watermarks during logits computation. Kirchenbauer et al. (2023) made minor modifications to current token logits based on hashes of previous k tokens (watermark logits). Lee et al. (2023) designed a similar watermarking method for low-entropy code generation. Wang et al. (2023) enabled multi-bit watermark information via a proxy LLM. Liu et al. (2023) implemented private watermarking through parameter sharing between the watermark generator and detector. In this work, we are more concerned with watermark robustness. Zhao et al. (2023) proved that using global fixed watermark logits (k=0) in Kirchenbauer et al. (2023)’s method achieves very strong attack robustness, yet Sadasivan et al. (2023)’s work shows that lower k values make watermarks vulnerable to security robustness (easy to break). Therefore, previous watermarking methods struggle to achieve both attack and security robustness. We propose a semantically invariant robust watermarking scheme to achieve decent attack and security robustness.
3 Preliminaries
We first introduce the necessary concepts used in this work. Language models could be divided into generative language models and embedding language models. A generative language model M takes a prompt $x^{prompt}$ and already generated text ${\bm{t}}_{:l-1}$ as input and generates the logits for the next token: $P_{M}(x^{prompt},{\bm{t}}_{:l-1})$. Meanwhile, an embedding language model $\mathrm{E}$ could generate an embedding $\mathrm{E}({\bm{t}})$ for the text ${\bm{t}}$. Usually, semantically similar texts will generate similar embeddings. For convenience in later sections, we use the language model to refer to generative language models.
A watermarking algorithm for large models embeds specific information in the generated text. The paradigm adopted in this paper is adding small watermark logits to the already generated next token logits. Specifically, the watermark logits can be defined as $P_{\mathrm{W}}({\bm{x}}^{prompt},\bm{t}_{:l-1})$, and the final logits could be defined as $P_{\hat{\mathrm{M}}}({\bm{x}}^{prompt},\bm{t}_{:l-1})=P_{\mathrm{M}}({\bm{x}}^{prompt},\bm{t}_{:l-1})+P_{\mathrm{W}}({\bm{x}}^{prompt},\bm{t}_{:l-1})$, where $\hat{\mathrm{M}}$ is the watermarked LLM. The watermark detector $P_{\mathrm{D}}$ corresponds to $P_{\mathrm{W}}$, outputting 1 if text ${\bm{t}}$ contains the watermark, otherwise 0. for consistency, we explicitly define $P_{\mathrm{W}}^{(i)}$ to denote the value at the i-th position of the watermark logits, and use $P_{\mathrm{W}_{i}}$ to denote a specific example of watermark logits.
The robustness of our work encompasses two aspects: attack robustness and security robustness. Attack robustness evaluates the probability that the watermarked text can still be correctly identified after semantically invariant modifications. Security robustness evaluates the accuracy of inferring the watermark rules from the watermarked text. Once the watermarking rules are disclosed, users can easily modify the text to remove the watermark, rendering attack robustness meaningless.
4 Proposed Method
In this section, we provide a detailed introduction to the proposed semantically invariant robust watermark algorithm named SIR. First, in Section 4.1, we introduce the overall process of watermark generation. Then, in Section 4.2, we explain the training process of the watermark model. Finally, in Section 4.3, we present the watermark detection process.
4.1 Watermark Generation
As discussed in the previous sections, one of the most important steps in the watermark generation process is the generation of the watermark logits. The watermark logits for the current token are usually determined by its preceding tokens. Our goal is to construct a continuous and robust mapping from the semantics of the preceding tokens to the watermark logits. In this way, semantic-preserving modifications to the text will only cause small perturbations to the watermark logits, thus ensuring the attack robustness of the entire watermark algorithm.
To extract semantic-invariant features of the text, we utilize an embedding language model $\mathrm{E}$ (e.g. BERT). Specifically, given a sequence of preceding tokens ${\bm{t}}_{:l-1}$, we first obtain its semantic embedding ${\bm{e}}_{l}=\mathrm{E}({\bm{t}}_{:l-1})$. To transform this semantic-invariant feature to a score over the vocabulary $\mathrm{V}$, we train a specialized watermark model $\mathrm{T}$ to generate the watermark logits: $P_{\mathrm{W}}=\mathrm{T}({\bm{e}}_{l})$. The overall algorithm is described in detail in Algorithm 1. A thorough introduction to the watermark model will be provided in Section 4.2. We will also provide a detailed analysis of the algorithm’s robustness in Section 5.
Algorithm 1 Watermark Generation
1:  Input: watermark strength $\delta$, a language model $\mathrm{M}$, previous generated text ${\bm{t}}=[t_{0}....t_{l-1}]$, a text embedding language model $\mathrm{E}$, a trained watermark model $\mathrm{T}$.
2:  Generate the next token logits from $P_{\mathrm{M}}$: $P_{\mathrm{M}}({\bm{x}}^{prompt},\bm{t}_{:l-1})$.
3:  Generate sentence embedding ${\bm{e}}_{l}=\mathrm{E}(\bm{t}_{:l-1})$.
4:  Generate watermark logit $P_{\mathrm{W}}$ from trained watermark model $\mathrm{T}({\bm{e}}_{l})$.
5:  Define a new language model $\hat{\mathrm{M}}$ where given input $\bm{t}=[t_{0}....t_{l-1}]$, the resulting logits satisfy
$P_{\hat{\mathrm{M}}}({\bm{x}}^{prompt},\bm{t}_{:l-1})=P_{\mathrm{M}}({\bm{x}}^{prompt},\bm{t}_{:l-1})+\delta\times P_{\mathrm{W}}({\bm{x}}^{prompt},\bm{t}_{:l-1}).$
6:  Output: watermarked next token logits $P_{\hat{\mathrm{M}}}(t_{l})$.
4.2 Watermark Model
In this section, we introduce the watermark model in detail. The goal of the watermark model is to convert semantic embeddings into watermark logits. First, we describe several desired properties for the similarity of watermark logits: semantic-consistent broad range, unbiased token preference, and balanced score.
To ensure the attack robustness, the similarity of the watermark logits should have a semantic-consistent broad range. That is, the similarity between generated watermark logits should be highly correlated with the similarity between text embeddings. Also, to ensure diversity and security, the similarty should have a broad range between 1 and -1 (language models typically generate embeddings with similarities concentrated between 0 and 1), as shown in the following equation:
$\forall x,y\in[-1,1],x<y,\exists i,j:\frac{P_{\mathrm{W}_{i}}\cdot P_{\mathrm{W}_{j}}}{||P_{\mathrm{W}_{i}}||_{2}\times||P_{\mathrm{W}_{j}}||_{2}}\in[x,y].$ (1)
Additionally, to further ensure the security robustness of the watermark, the watermark logits should have an unbiased token preference, meaning there should be no statistical preference for any token. This is shown in the following equation:
$\forall i\in\{1,2,\ldots,|\mathrm{V}|\},\sum_{j}P_{\mathrm{W}_{j}}^{(i)}=0.$ (2)
Moreover, to make the watermark stable and easy to detect, the watermark logits should have a balanced score, meaning the mean of the watermark logits is 0 and the score is uniform relative to the mean value, which is shown in the following equation:
$\forall j,\sum_{i=0}^{|\mathrm{V}|}\mathrm{sign}(P_{\mathrm{W}_{j}}^{(i)})=0,$ (3)
where $\mathrm{sign}(x)=1$ if $x$ is greater than 0, and -1 if $x$ is less than 0.
To achieve the above design goals, we devised a non-linear neural network (watermark model) to generate the watermark logits. This neural network consists of several fully connected layers with Rectified Linear Unit (ReLU) activations introduced to induce non-linearity. The watermark model employs two losses to meet the stated objectives: a similarity loss and a normalization loss.
The objective of the similarity loss is to attain a semantically consistent broad-range similarity for the watermark logits. Specifically, in order to achieve a broad range, the similarity of the text embeddings first needs to be normalized to $[-1,1]$ by calculating the mean of the similarity and then expanding the range using the $\mathrm{tanh}$ function. Finally, the overall goal of the similarity loss is to fit the similarity between the watermark logits to the transformed text embedding similarities. The standard definition of the similarity loss $\mathcal{L}_{s}$ is as follows:
$\sum_{i}\sum_{j}|\frac{\mathrm{T}({\bm{e}}_{i})\cdot\mathrm{T}({\bm{e}}_{j})}{||\mathrm{T}({\bm{e}}_{i})||_{2}\times||\mathrm{T}({\bm{e}}_{j})||_{2}}-\mathrm{tanh}(k_{1}(\frac{{\bm{e}}_{i}\cdot{\bm{e}}_{j}}{||{\bm{e}}_{i}||_{2}\times||{\bm{e}}_{j}||_{2}}-\sum_{k}\sum_{l}\frac{{\bm{e}}_{k}\cdot{\bm{e}}_{l}}{|N|^{2}||{\bm{e}}_{k}||_{2}\times||{\bm{e}}_{l}||_{2}}))|,$ (4)
where $|N|$ is the size of the sentence embedding space, $\mathrm{T}$ is the watermark model, and $k_{1}$ is a hyperparameter used to adjust the range of similarity.
We then introduce the normalization loss, whose goal is to have the watermark logits exhibit an unbiased token preference while maintaining a balanced score. Specifically, this requires the mean of each watermark logit to be zero, and the mean over all tokens in the watermark logits to also be zero. Finally, to make the score of the watermark logits uniform, we constrain the absolute value of each value to be close, with the standard definition of the normalization loss as follows:
$\mathcal{L}_{n}=\sum_{i}|\sum_{j}\mathrm{T}({\bm{e}}_{i})^{(j)}|+\sum_{i}|\sum_{j}\mathrm{T}({\bm{e}}_{j})^{(i)}|+\lambda_{1}\sum_{i}\sum_{j}|R-\mathrm{T}({\bm{e}}_{j})^{(i)}|,$ (5)
Where R is a hyperparameter denoting the target absolute value for each value in the watermark logits, and $\lambda_{1}$ is a weight conditioning the normalization loss internally.
The final training loss is a weighted sum of the similarity loss and normalization loss, expressed as follows, where $\lambda_{2}$ is a hyperparameter balancing the two losses:
$\mathcal{L}=\mathcal{L}_{s}+\lambda_{2}\mathcal{L}_{n}.$ (6)
To augment the separability of watermark logits, the output from the watermark model is further processed by the $\mathrm{tanh}$ function, yielding the final watermark logits, denoted as: $\mathrm{tanh}(k_{2}\,T(\mathbf{e}_{i}))$.
4.3 Watermark Detection
In this section, we introduce the method for detecting the watermark.
Analogous to watermark generation, the watermark logit is first computed per token via Algorithm 1. The score at the position of each token $t_{j}$ is selected. Averaging all scores gives the detection score:
$z=\frac{\sum_{j=1}^{N}P_{\mathrm{W}}^{(j)}({\bm{x}}^{prompt},\bm{t}_{:j-1})}{N}.$ (7)
Without a watermark, the expected score is 0 since the watermark logit mean is 0. When a watermark is present, the score substantially exceeds 0, making this a suitable validation approach. We utilize the z-score in the following to indicate our detection score.
5 Robustness Analysis
This section analyzes the attack robustness of the watermark algorithm. As detection relies on the z-score, robustness can be assessed by the z-score change with modifications to the generated text ${\bm{t}}_{:N}$. The z-score change magnitude indicates robustness. We use ${\bm{t}}^{\prime}_{:N}$ to denote the modified text.
Let $U$ represent the set of altered tokens Their new scores are considered zero. For unmodified tokens, score changes result from embedding alterations of the proceeding tokens:
$|\Delta z|\leq\frac{\sum_{j\in U}|P_{\mathrm{W}}({\bm{x}}^{prompt},\bm{t}_{:j-1})|+\sum_{j\notin U}|P_{\mathrm{W}}({\bm{x}}^{prompt},\bm{t}_{:j-1})-P_{\mathrm{W}}({\bm{x}}^{prompt},\bm{t}^{\prime}_{:j-1})|}{N}.$ (8)
Since watermark logit generation from text embeddings is continuous, the Lipschitz constant L can bound the inequality per properties of continuous functions:
$|\Delta z|\leq\frac{\sum_{j\in U}|P_{\mathrm{W}}({\bm{x}}^{prompt},\bm{t}_{:j-1})|+\sum_{j\notin U}L|\mathrm{E}(\bm{t}_{:j-1})-\mathrm{E}(\bm{t}^{\prime}_{:j-1})|}{N}.$ (9)
This shows the watermark is theoretically robust as long as text modifications do not drastically alter semantics, such that embeddings stay similar. In fact: significant semantic changes would make the watermark meaningless. Moreover, an analysis of security robustness is provided in the appendix.
6 Experiment
6.1 Experiment settings
Dataset and Prompt: Similar to the previous works Kirchenbauer et al. (2023), we utilize the C4 dataset (Raffel et al., 2020) for data generation, taking the first 30 tokens as prompts and generating the next 200 tokens. The original C4 texts serve as human-written examples. The test objective is to distinguish between generated text and human-written text. During training of the watermark model, we utilize the WikiText-103 dataset (Merity et al., 2016) (different from C4) to generate embeddings.
Baseline and Language Model: We selected two watermarking algorithms as baselines. One is KGW-k (Kirchenbauer et al., 2023) where k is the number of preceding tokens to hash. The other is exponential minimum sampling (EXP-edit) (Kuditipudi et al., 2023), an approach that introduces watermarks through a pre-selected sequence of sampling probabilities during the sampling process and adopts edit distance-based detection to achieve strong robustness. For language models, we use LLaMA-7B (Touvron et al., 2023), OPT1.3B, and OPT2.7B (Zhang et al., 2022) for text generation. Additionally, we tested the efficacy of the watermarking algorithm under both stochastic and deterministic scenarios using sampling and beam search decoding algorithms, respectively. For embedding language models, we utilized Compositional-BERT (Chanchani & Huang, 2023) due to its superior ability to produce embeddings that better distinguish text.
Evaluation: Similar to Zhao et al. (2023)’s method, to avoid the impact of detection thresholds, we set false positive rates at 1% and 10% and adjusted the detector’s thresholds accordingly. For comparison, we also report F1 scores at optimal thresholds. Further, we use the superior LLaMA-13B model for perplexity evaluation. For safety robustness assessment, we adopt Sadasivan et al. (2023)’s approach of analyzing occurrence frequencies of 181 common words. See Section 6.3 for details.
Hyper-parameters: The watermark model uses a four-layer fully connected residual network with rectified linear unit activations. Hyperparameters are set to $k_{1}=20$, $k_{2}=1000$, $\lambda_{1}=10$, $\lambda_{2}=0.1$, and the Adam optimizer (lr=1e-5) is used for training. The detailed network architecture is provided in the appendix. All experiments were conducted using the NVIDIA Tesla V100 32G GPU.
6.2 Attack Robustness Analysis
In Table 1, we list the detection accuracy of our watermark algorithm and various baseline algorithms under the no-attack setting and when texts are rewritten using GPT3.5 as well as two different DIPPER settings (Krishna et al., 2023). For GPT3.5, we use the gpt-3.5-turbo-0613 version with the prompt Rewrite the following paragraph:. For DIPPER-1 the lex diversity is 60 without order diversity, and for DIPPER-2 we additionally increase the order diversity by 20. EXP-Edit cannot be used under beam search so we do not list its result.
Table 1 shows that our watermarking algorithm achieves strong robustness against all rewrite attacks. Specifically, for watermarked texts rewritten by GPT-3.5 and two DIPPER rewriters, our algorithm still achieves an average detection F1 score of 0.93. This performance is comparable to that of the KGW-1 algorithm, markedly surpassing other watermark algorithms, including KGW-2, KGW-4, and EXP-Edit. As modifications to the token in the KGW-1 algorithm only affect the label of that particular token, this represents the upper bound of attack robustness for watermark algorithms based on watermark logits. Nonetheless, subsequent experiments demonstrate that the KGW-1 algorithm possesses poor security robustness.
To further demonstrate the robustness of our watermarking method against semantic-preserving text attacks, Figure 2(a) compares the detection F1 score of our watermark algorithm and other baseline methods under different synonymous word substitution scenarios. The synonyms are obtained from the WordNet synset (Miller, 1995). Since replacing words solely could still alter the semantics as many synonyms are only applicable in certain contexts, we additionally employ BERT (Devlin et al., 2018) to constrain the synonym substitution to induce minimal embedding changes as a context replacement approach. Without context replacement, our algorithm demonstrates robustness comparable to KGW-2, while under semantics-preserving substitutions, our method achieves near state-of-the-art robustness against attacks, similar to KGW-1 and EXP-Edit. These results clearly demonstrate the robustness of our method for semantically invariant text modification.
6.3 Security Robustness Analysis
Table 1: We compared the performance of our watermarking method with others, including KGW-k (Kirchenbauer et al., 2023) and EXP (Kuditipudi et al., 2023), using text generated by LLaMA-7B. Tests involved watermark detection accuracy under no attack, GPT3.5 rewrite attacks, and two DIPPER (Krishna et al., 2023) settings.
Sampling Beam search
Setting Method 1% FPR 10% FPR Best 1% FPR 10% FPR Best
TPR F1 TPR F1 F1 TPR F1 TPR F1 F1
No attack KGW-1 0.960 0.975 1.000 0.952 0.983 1.000 0.995 1.000 0.952 0.998
KGW-2 1.000 0.995 1.000 0.952 0.998 1.000 0.995 1.000 0.952 1.000
KGW-4 1.000 0.995 1.000 0.952 0.998 1.000 0.995 1.000 0.952 1.000
EXP-Edit 0.995 0.993 1.000 0.952 0.995 $\times$ $\times$ $\times$ $\times$ $\times$
SIR(ours) 1.000 0.995 1.000 0.952 0.995 1.000 0.995 1.000 0.952 1.000
GPT3.5 KGW-1 0.590 0.738 0.885 0.891 0.905 0.890 0.937 0.965 0.935 0.955
KGW-2 0.535 0.693 0.760 0.817 0.823 0.655 0.787 0.795 0.839 0.865
KGW-4 0.225 0.364 0.490 0.614 0.705 0.420 0.587 0.660 0.750 0.795
EXP-Edit 0.435 0.602 0.645 0.739 0.775 $\times$ $\times$ $\times$ $\times$ $\times$
SIR(ours) 0.740 0.856 0.865 0.880 0.900 0.805 0.887 0.945 0.924 0.938
DIPPER-1 KGW-1 0.715 0.829 0.940 0.922 0.930 0.930 0.959 0.975 0.939 0.962
KGW-2 0.450 0.616 0.710 0.785 0.815 0.770 0.865 0.880 0.888 0.908
KGW-4 0.220 0.358 0.545 0.627 0.728 0.380 0.547 0.765 0.820 0.843
EXP-Edit 0.630 0.768 0.740 0.804 0.830 $\times$ $\times$ $\times$ $\times$ $\times$
SIR(ours) 0.765 0.862 0.905 0.903 0.920 0.890 0.937 0.950 0.927 0.948
DIPPER-2 KGW-1 0.765 0.862 0.935 0.918 0.925 0.910 0.948 0.975 0.940 0.960
KGW-2 0.470 0.635 0.685 0.768 0.803 0.725 0.838 0.860 0.878 0.898
KGW-4 0.150 0.259 0.475 0.603 0.718 0.315 0.475 0.645 0.739 0.783
EXP-Edit 0.485 0.649 0.635 0.732 0.775 $\times$ $\times$ $\times$ $\times$ $\times$
SIR(ours) 0.875 0.928 0.92 0.911 0.931 0.870 0.926 0.940 0.922 0.940
In Figure 3(a), we analyze the security robustness of our method. Security robustness refers to the difficulty of cracking the watermarking rules. In this work, we adopt the spoofing attack method of Sadasivan et al. (2023), which analyzes the word frequencies of 181 commonly occurring words. If a word in watermarked text has a higher frequency than in natural text, it is deemed watermarked (with positive probability in the watermark logits). The accuracy of this identification is called watermark decryption accuracy, which measures the security robustness of the watermark.
Specifically, for KGW-k we count the frequency of the last word given fixed k-1 prefixes. Due to the semantic relevance of our watermarking rules, we employed DBpedia Class dataset (Gangemi et al., 2012) to examine watermark decryption accuracy. Specifically, we analyzed the accuracy across three different levels: the overall class, L1 class (e.g., species), and L2 class (e.g., species: animal). Due to insufficient data at the L3 class level, our research focuses solely on L1 and L2 classes.
Figure 3(a) illustrates the trade-off between attack robustness (measured by detection accuracy after rewrite) and security robustness (measured by watermark decryption accuracy) in Kirchenbauer et al. (2023)’s watermarking algorithm. A smaller k implies stronger attack robustness yet simpler watermarking rules and thus lower security robustness. Our watermarking algorithm achieves attack robustness close to KGW-1 while maintaining security robustness close to KGW-4. Although our security robustness degrades slightly with more refined domain text (L1 and L2 level classes), even for the very refined L2 class it is still far superior to KGW-3. Therefore, our watermarking algorithm successfully addresses the attack robustness versus security robustness trade-off in prior work.
Refer to captionRefer to caption
Figure 2: The left figure shows how detection accuracy changes for different watermark models as the synonym replacement ratio increases. The middle figure shows the correlation between embedding similarity generated by the embedding model and the similarity of the generated watermark logits. The right figure illustrates watermark logits with and without the normalization loss.
Figure 3: The left figure depicts the trade-off between security robustness (measured by watermark decryption accuracy) and attack robustness (measured by F1 score after rewrite) across different watermarking algorithms. The right figure shows the text quality generated by language models with different watermarking methods (measured by text perplexity).
Table 2: Text generation speed, measured in seconds (s), was compared with and without applying the watermarking algorithm, for a generated token length of 200. All experiments were performed on a single NVIDIA Tesla V100 32GB GPU.
Embedding Model Setting OPT-1.3B OPT-2.7B LLaMA-7B
Com-BERT Base w/o watermark 3.14 4.10 5.76
w/ watermark 4.98 6.87 7.23
w/ watermark (parallel) 3.32 4.45 5.81
Com-BERT Large w/o watermark 3.14 4.10 5.76
w/ watermark 5.67 7.44 7.83
w/ watermark (parallel) 3.55 4.89 5.94
6.4 Watermark Model Analysis
To better analyze our watermarking algorithm, Figure 2(b) shows the relationship between the similarity of text embeddings and the similarity of watermark logits generated by the watermark model. The average similarity of watermark logits is shown for each range of embedding similarity. It can be seen that the similarity of watermark logits maintains a strong correlation with the embedding similarity, especially for very similar embeddings that certainly correspond to very similar watermark logits (similarity greater than 0.8). Meanwhile, the range of watermark logit similarity is also extended to -1 to 1 (the figure only shows the average similarity in the range so -1 is not shown), which also makes the watermarking rules more diverse and robust. From the figure, the average similarity of the original sentence embeddings is 0.45, which also reflects the advantage of using Compositional-BERT, which can better distinguish dissimilar texts compared to the original BERT, helping to improve the diversity of watermarks. We will analyze in more detail in the appendix the impact of different embedding models on watermarking. Figure 2(b) also shows our normalization loss results in less distinguished watermark logits.
In Figure 2(c), we plot the specific watermark logits. It shows that after using our normalization loss, the watermark logits becomes more symmetrical without overly small logit values. After applying the tanh scaling method introduced in Section 4.3, the watermark logits only contain values close to -1 and 1. The potential impact of the watermark on text quality depends on the maximum absolute token logit value (as it would only select that token if too large), so making the absolute values of all token logits equal to the minimum max value is an optimal solution. See more detail in the Appendix.
6.5 Time Complexity and Text Quality
In Table 2, we further analyzed the time complexity of our watermarking algorithm. As shown, without parallel processing, our algorithm imposes additional time costs on text generation, mainly due to the language model embedding and the watermark model is just a four-layer linear network, with negligible time costs. As language models increase in size, the proportional time increase diminishes, with LLaMA-7B only taking 40% longer. Figure 1 illustrates that our algorithm can generate LLM logits and watermark logits in parallel with minimal additional time cost, particularly for larger models like LLaMA-7B.
To evaluate the impact of our watermarking method on text quality, Figure 3(b) shows the perplexity comparison of the generated text. The perplexity calculation uses the LLaMA-13B model, and the text on OPT2.7B and LLaMA-7B is generated using sampling decoding. It can be seen that our method and previous watermarking methods have a slight impact on text quality (perplexity increases slightly). But at the same time, our method can achieve slightly lower perplexity than the KGW series of watermarking algorithms, which may be the result of our semantic-based watermarking algorithm being more consistent in text expression. We will present examples in the appendix.
7 Conclusion
In this work, we propose a semantically invariant robust watermarking algorithm that generates watermark logits based on the semantics of context. We validate the robustness of our algorithm through a series of analyses and experiments, particularly in scenarios involving text rewriting and synonym substitution, while maintaining the watermark’s resistance to decryption. For future improvements, we suggest utilizing superior quality embedding models, which will enhance the performance of our watermarking algorithms. Furthermore, there is potential for expansion of our algorithm for use in different watermarking approaches, such as employing pre-determined probability sequences for token sampling.
Appendix A Case Study
To more intuitively demonstrate the robustness of our semantic invariant watermarking algorithm, we present a comparison between texts generated by our algorithm and KGW-1, KGW-2, and KGW-4 given the same prompt in Figures 4 and 5. Both figures show the original texts produced by the corresponding watermark algorithms and the texts rewritten by GPT-3.5. The green color indicates tokens with watermark logit values greater than 0 (corresponding to the green tokens in the KGW-k algorithms). These examples illustrate that our algorithm maintains high z-scores even after GPT-3.5 rewriting, while the robustness of the KGW-k algorithms decreases as k increases.
Figure 5: This figure demonstrates the examples of the KGW-2 method and the KGW-4 method when using the same prompt. The other settings of this figure are identical to Figure 4
Appendix B Watermark Logits Analysis
In this section, we analyze the influence of the shape of watermark logits on watermarks. Specifically, we study the impact on watermark detection success rate and text quality after four different transformations of the original watermark logits. The four transformations are:
  • $\mathrm{tanh}(1000\*{\bm{x}})$ scaling, which is the method used in this work, after which all values will be close to 1 or -1, as shown in the top graph of Figure 6(a).
  • Linear scaling: uniformly distributed between -1 and 1 according to the rank of the data (top of Figure 6(b)):
    $L({\bm{x}})=-1+2\times\frac{\mathrm{argsort}(\mathrm{argsort}(\mathbf{{\bm{x}}}\text{))}}{\mathrm{len}(\mathbf{{\bm{x}}})-1}$ (10)
  • $\mathrm{tanh}(10\*L({\bm{x}}))$ scaling, applying an additional $\mathrm{tanh}$ transform on top of the linear scaling, as shown in the bottom graph of Figure 6(a).
  • $L(x)^{3}$ scaling, applying an additional cubic transform on top of the linear scaling, as shown in the bottom of Figure 6(b).
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 6: Examples of the shapes of four watermark logits that we tested are presented: (a) The top graph of Figure (a) illustrates the $\mathrm{tanh(1000x)}$ utilized, (b) the top graph of Figure (b) displays the data originally ranked and linearly scaled to the interval between -1 and 1, as demonstrated by Equation 10. Subsequently, the graph below Figure (a) depicts a $\mathrm{tanh}$ transformation applied to the top graph of Figure (b), and the graph below Figure (b) demonstrates a cubic transformation applied likewise to the top graph of Figure (b).
Figure 6 illustrates that the four distinct watermark logits exhibit minimal differences in their impact on text perplexity (PPL) under various $\delta$ values, corroborating our hypothesis articulated in Section 6.4. This hypothesis posits that the influence on text quality predominantly depends on the token with the maximum value in the watermark logits, as the four employed watermark logits possess the same maximum value under identical $\delta$ values. Nonetheless, these watermark logits require different $\delta$ values to achieve optimal detection results; the need for larger $\delta$ values increases with a higher distribution of values near zero. Correspondingly, the watermark logits we adopt can achieve excellent detection performance with minimal impact on text quality.
Appendix C Analysis of Different Embedding Models
To investigate the impact of different embedding language models on watermark algorithms and to discern how to appropriately select an embedding model, we enumerate in Figure 8 the distribution of sentence embeddings and the relationship between sentence embedding similarities and watermark logits similarities when employing three distinct embedding models: BERT Devlin et al. (2018), Sentence-BERT Reimers & Gurevych (2020), and Compositional-BERT Chanchani & Huang (2023). Sentence-BERT (SBERT) and Compositional-BERT (CBERT) have been refined for sentence similarity tasks. They show a near-normal distribution of embedding similarities with moderate average similarity, making them ideal as our embedding models. In this study, we opt for Compositional-BERT, though employing Sentence-BERT would yield comparable results. Conversely, the origin BERT model, not specifically designed for text similarity, tends to have uniformly high text similarity scores, which in turn means that minor perturbations in embeddings can significantly affect watermark logits, rendering it unsuitable as the embedding language model for our robust watermark algorithm.
Figure 7: The left figure shows how the watermark detection F1 score changes as the value of $\delta$ (the extent of watermark augmentation) varies, using the 4 different watermark logits from Figure 6. The right figure shows how the perplexity (PPL) of the generated text changes with $\delta$ for the 4 different watermark logits.
Figure 8: Comparison of similarity distributions for different embedding language models and the corresponding watermark logits after transformation.
Appendix D Detail Network Structure of the Watermark Model
To elaborate in detail on our methodology, we present the Python implementation code of the watermark model described in this work, as illustrated in Figure 9. Precisely, the network consists of an input layer, output layer, and a middle linear network with residual connections. The dimensions of the input correspond with the output dimensions of the embedding model, while the output dimensions align with the size of the vocabulary. During our implementation, given the variation in vocabulary size across different language models, we predetermined a fixed output dimension of 1000. Subsequently, a random mapping was established to project various vocabularies randomly onto this output dimension.
Furthermore, it is unnecessary to recalculate using the watermark model for every token generated during the implementation process. Since the semantics of the text typically do not alter dramatically with the addition of individual tokens, we compute the watermark logits at intervals of N steps (where N ranges between 5 and 10) in practice. This approach effectively reduces the computational complexity of our watermarking algorithm.
1class ResidualBlock(nn.Module):
2    def __init__(self, dim):
3        super(ResidualBlock, self).__init__()
4        self.fc = nn.Linear(dim, dim)
5        self.relu = nn.ReLU()
6
7    def forward(self, x):
8        out = self.fc(x)
9        out = self.relu(out)
10        out = out + x
11        return out
12
13class TransformModel(nn.Module):
14    def __init__(self, num_layers, input_dim, hidden_dim, output_dim):
15        super(TransformModel, self).__init__()
16        self.layers = nn.ModuleList()
17        self.layers.append(nn.Linear(input_dim, hidden_dim))
18        for _ in range(num_layers - 2):
19            self.layers.append(ResidualBlock(hidden_dim))
20        self.layers.append(nn.Linear(hidden_dim, output_dim))
21
22    def forward(self, x):
23        for i in range(len(self.layers)):
24            x = self.layers[i](x)
25        return x
Figure 9: The code implementation of the watermark model in this paper, where the TransformModel class represents the watermark model network.
Appendix E SECURITY ROBUSTNESS ANALYSIS
In this section, we provide a general analysis of the security robustness. Herein, security robustness refers to the likelihood of users deducing the watermark generation methodology from knowledge of the watermark algorithm and generated watermark text examples.
More specifically, there is a strong correlation between security robustness and the number of watermark rules in the watermark algorithms. For instance, the KGW-k watermark algorithm possesses a number of watermark rules equivalent to $|V|^{k}$ (each combination of tokens of $k$ window size corresponds to a watermark logits), hence exhibiting an exponential increase in security robustness with increasing $k$.
Regarding our algorithm, assuming no correlation between different watermark generation rules, the number of rules in our algorithm can be calculated using the following equation:
$\sum_{i=1}^{T}|V|^{i}$ (11)
where $T$ represents potential text length. However, many rules generated in this manner are semantically similar, necessitating consideration of rule correlation when estimating the number of watermark rules. Calculating such correlation is highly complex and not straightforward.
Rather than directly calculating the complexity of our watermarking scheme, we demonstrate its security and robustness by the success rate of attack algorithms in cracking the watermarking rules. Thus in Figure 3(a) we utilize the word frequency analysis method employed by Sadasivan et al. (2023). In this experimental setup, we analyzed 2000 generated texts, each with a length of 200, and found that the security robustness of our algorithm is superior to KGW-3 and closely approximates KGW-4.